what is the equation of directors circle in hyperbola and ellipse

Grade:12

2 Answers

naveenkumar arepally
8 years ago
An ellipse, its minimum bounding box, and its director circle. In geometry, the director circle of an ellipse or hyperbola (also called the orthoptic circle or Fermat–Apollonius circle) is a circle formed by the points where two perpendicular tangent lines to the curve cross.The length of latus rectum of (i) Parabola is 4a (ii) Ellipse is 2b2a (iii) Hyperbola is ... Equations of director circle of— y2b2 = 1 is x2 + y2 = a2 + b2 .The length of latus rectum in (i) parabola is 4a; (ii) ellipse and hyperbola is 2bVa. 5. ... The equation of the director circle is Ii) Ellipse is : x2 + y2 = a2 + b2
SHANMUKESHWAR
8 years ago
Equation of the director circle of the hyperbola x2/a2 – y2/b2 = 1 is x2 + y2 = a2 – b2
the equation of its auxiliary circle of the ellipse x2 + y2 = a2 .